Calculate the pH and the pOH of the solution at 25 c for 0.0031 M Ca(OH)2


1
Expert's answer
2021-03-01T02:37:23-0500

Solution:

When Ca(OH)2 molecules dissolve they dissociate into Ca2+ ions and OH- ions.

Ca(OH)2 → Ca2+ + 2OH-

For every Ca(OH)you produce 2OH-.

Therefore, the [OH-] equals 0.0062 M:

[OH-] = 2 × CM(Ca(OH)2) = 2 × 0.0031 M = 0.0062 M


We can convert between [OH-] and pOH using the following equations:

pOH = - log[OH-]

pOH = -log(0.0062) = 2.21

pOH = 2.21


For any aqueous solution at 25C:

pH + pOH = 14

pH = 14 - pOH = 14 - 2.21 = 11.79

pH = 11.79


Answer: pOH = 2.21; pH = 11.79